# Nanoscale magnetometry of a synthetic three-dimensional spin texture

### v1

@@ Abstract
Multilayered synthetic antiferromagnets (SAFs) are artificial three-dimensional (3D) architectures engineered to create novel, complex, and stable spin textures. Non-invasive and quantitative nanoscale magnetic imaging of the two-dimensional stray field profile at the sample surface is essential for understanding the fundamental properties of the spin-structure and being able to tailor them to achieve new functionalities. However, the deterministic detection of spin textures and their quantitative characterization on the nanoscale remain challenging. Here, we use nitrogen-vacancy scanning probe microscopy (NV-SPM) under ambient conditions to perform the first quantitative vector-field magnetometry measurements in the multilayered SAF [(Co/Pt)5/Co/Ru]3/(Co/Pt)6. We investigate nanoscale static and dynamic properties of antiferromagnetic domains with boundaries hosting "one-dimensional" ferromagnetic stripes with 100 nm of width and periodic modulation of the magnetization. By employing NV-SPM measurements in different imaging modes and involving NV-probes with various crystallographic orientations, we demonstrated distinct fingerprints emerging from GHz-range spin noise and constant stray fields on the order of several mT. This provides quantitative insights into the structure of domains and domain walls, as well as, into magnetic noise associated with thermal spin-waves. Our work opens up new opportunities for quantitative vector-field magnetometry of modern magnetic materials with tailored 3D spin textures and stray field profiles, and potentially novel spin-wave dispersions—in a quantitative and non-invasive manner, with exceptional magnetic sensitivity and nanometer scale spatial resolution.
@@ 1 Introduction
Synthetic antiferromagnets (SAFs) are artificial magnetic systems that offer unique opportunities for fundamental research and applications due to their highly tunable static and dynamic properties Duine2018, Wang2023, Wang2024-bk, Xu2021, Rong2024. In particular, SAFs present stable magnetic textures and high-frequency spin dynamics, which are ideal properties for magnetic storage technologies and spintronic devices Legrand2020, Pham2024, Godinho2024, Chen2021, Lonsky2022. Multilayered SAFs are systems with a more complex layered structure, designed to tune magnetic energy terms to create novel magnetic textures. Hellwig2003, HELLWIG2007, Böhm2022, Mouhoub2023, Wang2023. For instance, systems with perpendicular magnetic anisotropy (PMA) can be engineered to stabilize stripe and bubble domain states Kiselev2010, to induce a domain wall (DW) transition from antiferromagnetic (AF) DWs to DWs with ferromagnetic (FM) coresHellwig2003-prl, or to stabilize FM domains separated by AF DWs when the interlayer exchange energy is reduced below the demagnetization energy Salikhov2025. From the dynamic point of view, dipolar interaction can affect the spin-wave modes and induce nonreciprocity on the spin-wave dispersion Gallardo2019, Jimenez-Bustamante2025-ub, giving rise to complex spin-wave propagation. Additionally, multilayered structures with periodic magnetic stripe domain patterns exhibit spin-wave modes and dispersion relations analogous to those of a one-dimensional magnonic crystal, which can be used to control spin-wave propagation.Banerjee2017, GRUSZECKI202129. Therefore, these types of multilayered architectures are a powerful and application friendly approach to develop new magnetic materials with spin textures and properties useful to improve the performance of SAFs and to design the next generation of magnetic devices Gubbiotti_2025, Zhang2021, Lavrijsen2013, Fernandey2016.
@@
The direct imaging of the two-dimensional stray magnetic field profile at the sample surface with nanoscale spatial resolution can provide insights into the properties of individual domains and DWs, which is a crucial step toward understanding and improving the fundamental properties and performance of multilayered SAFs. Magnetic force microscopy (MFM) combined with macroscopic magnetometry have been employed to investigate multilayered SAFs and to connect the nanoscale magnetism of every domain/DW with the macroscopically observed properties, such as magnetization or magnetic hysteresis Hellwig2003, Hellwig2003-prl, HELLWIG2007, Salikhov2025, Böhm2022, Koch2020, Salikhov2021, Salikhov2022. However, the field of a few tens of mT of the MFM probe Balasubramanian2008-ec can induce reversible changes or irreversible reorientation of the magnetic moments while scanning the sample Stepanova2022, Zelent2021, Casiraghi2019-wd, hindering access to the intrinsic undisturbed magnetic texture. Furthermore, macroscopic magnetometry provides only information about the average behavior of a large number of magnetic domains. Moreover, many magnetic imaging techniques probe only static properties non-quantitatively and lack the sensitivity required to investigate phenomena with weak stray magnetic field (such as AF order). Therefore, despite the enormous effort in the last few years, there is still a remaining challenge regarding the deterministic detection of magnetic textures in multilayered SAFs, along with the local, quantitative, and non-invasive characterization of both static and dynamic properties.
@@
Nitrogen-Vacancy scanning probe microscopy (NV-SPM) has recently gained immense interest as a robust quantum sensing tool. It utilizes the quantum properties of a single NV defect in diamond for quantitative and non-invasive magnetic imaging, where the magnetic vector-field sensitivity can be as high as nT with an exceptional nanoscale spatial resolution, in the range of tens of nanometers Degen2017, Finco2023, Rovny2024-bz, Budakian_2024, Christensen_2024, Degen2025. NV-SPM has been previously employed to investigate ultra-thin SAFs or FM thin films to determine the nature of DWs and strength of the Dzyaloshinskii-Moriya interaction Tetienne2014, Tetienne2015-rc, Gross2016, DW motion Tetienne-2014, McLaughlin2023, spin-wave dispersion and magnetic noise Finco2021-ne, Finco2025. In such systems, the magnetic thickness is typically less than 2 nm with a few mT of stray field, for which unwanted effects that often make quantitative measurements difficult, such as spin-mixing of the NV-spin states and suppression of the quantum sensor’s signal due to high off-axis stray fields, are not present. Consequently, magnetic imaging of thick multilayered SAFs with complex and novel spin textures and dynamics, producing both a large stray field of tens of mT and spin noise, remained unexplored using NV-SPM.
@@
In this work, we use NV-SPM under ambient conditions to investigate the magnetic texture of the multilayered SAF of the type [(Co/Pt)X-1/Co/Ru]N-1/(Co/Pt)X with , and . We explore the domains and DWs structure, as well as their static and dynamic properties, by performing measurements in different imaging modes and employing NV-probes with various crystallographic orientations to identify distinctive fingerprints arising from both constant stray magnetic fields and spin noise. We use optically detected magnetic resonance (ODMR) and relaxometry measurements to demonstrate the presence of both sample stray fields on the order of mT and spin noise in the GHz range. We perform qualitative magnetic imaging with commercial diamond probes containing a single NV and fabricated from (100)- and (110)-oriented diamond crystalsWelter_2022. By using a (111)-oriented single NV probeRohner_2019, we quantitatively resolve the nanoscale DW structure and the spatial distribution of the sample stray field. Combining this with micromagnetic simulation, we confirm a three-dimensional (3D) model of the spin texture that agrees well with our quantitative measurements. Our work opens novel opportunities to understand and characterize static and dynamic properties in layered architectures with complex magnetic textures and spin-waves, quantitatively and non-invasively, with exceptionally high magnetic sensitivity and nanometer-scale spatial resolution.
@@ 2.1 NV-SPM imaging in a multilayered SAF
As illustrated in Figure 1(a), NV-SPM employs a single NV defect in a diamond scanning probe (see section 1 in supplementary information (SI)). The NV sensor is used to map the sample surface and spatially resolve the magnetic texture. Due to its spin-dependent photoluminescence (PL), the NV-spin state can be optically initialized with a green laser (515 nm) and read out by detecting the red NV-PL signal in the wavelength range of 650-850 nm (see methods). Qualitative magnetic imaging can be performed by mapping the intensity of the NV-PL signal while scanning the sample. For direct quantification of the sample stray field by ODMR, microwave (MW) excitation is applied to promote the NV spin transition. Thereby, the stray magnetic field from the sample is determined from the Zeeman splitting produced by its component along the NV-axis, which is oriented along the ⟨111⟩ crystallographic direction in the diamond probe. The orientation of the NV-axis is defined by the angles in the laboratory frame. From our ODMR measurements under ambient conditions, we estimate a DC magnetic field sensitivity of 2 T/. The value of the detected stray field and the spatial resolution for magnetic imaging depend on the NV-sample distance, which is 36-76 nm for all our measurements (see Section 2 in SI).
@@
Figure 1(b) illustrates the structure of the multilayered SAF investigated in this work. The stacking of 6 Co layers with 5 Pt interlayers give rise to FM building blocks with PMA. i.e. out-of-plane (OOP) magnetization parallel or antiparallel to the film normal axis. The insertion of a non-magnetic material like Ru introduces an AF-coupling between adjacent FM blocks through the Ruderman–Kittel–Kasuya–Yosida exchange interaction. The strength of the AF-exchange interaction can be tuned via the Ru thickness. Duine2018, Wang2023, HELLWIG2007. In our sample, we have the AF-coupling of four FM blocks, and the full structure is described as [[Co(0.55 nm)/Pt(0.70 nm)]5/Co(0.55 nm)/Ru(0.75 nm)]3/[Co(0.55 nm)/Pt(0.70 nm)]6. The numbers in nm correspond to the thickness of every thin film along the stacking. Thus the magnetic thickness of a single FM block is nm. The hysteresis curve measured on this sample shows four discrete steps (see section 3 in SI), where every step corresponds to the magnetic reversal of one of the four FM blocks, with an average saturation magnetization kA/m for configurations with a full magnetization pointing up/down.
@@
MFM images recorded after sample saturation with an in-plane applied field reveal a remanent state constituted by one-dimensional FM stripe domains with periodic bright and dark contrast at the boundary between AF domainsHellwig2003-prl, HELLWIG2007, as shown in Figure 1(c). The AF domain sizes are primarily in the micrometer range, with 100 nm-wide FM stripes and average periodicity of the contrast of 300 nm (see section 3 in SI) . The width of the stripes and the periodicity of the contrast can vary slightly from one stripe to another, but typically remain on the order of hundreds of nanometers. The alternating bright and dark contrast observed along the FM stripe, as well as, the slight difference in magnetic contrast between adjacent AF domains, originates from the parallel and antiparallel alignment of the sample stray field and the magnetic MFM tip. Such magnetic texture has been observed by Hellwig et al. in similar samples with various , , and layer thicknesses Hellwig2003-prl, HELLWIG2007. However, gaining further quantitative microscopic details on the spin texture of domains and DWs along with their dynamic properties are beyond the scope of standard MFM measurements.
@@
To this end, we employ the NV-SPM approach. Figure 1(d) shows an image acquired in the NV-PL quenching mode with a (100)-oriented NV probe. The NV-PL signal is detected pixel by pixel as the sample is scanned across the NV tip which is kept stationary at the focus of the initializing green laser beam. MW excitation is not applied during the measurements in this imaging mode. We observe large bright areas on the AF domains. Along the FM stripe domains the PL signal is quenched by about 40%. The NV-PL signal can be quenched either due to spin-mixing effects induced by strong off-axis stray fields Tetienne_2012 or due to a decrease of the spin-relaxation time () in case spin noise is present with frequencies resonant with the NV spin transition Finco2021-ne, Rollo2021. Contrary to MFM measurements, there is no periodic variation of the contrast. However, MFM-like contrast can be achieved in NV-PL maps only after applying a suitable off-axis (not aligned with the NV-axis) external magnetic field, as described in the following section.
@@ 2.2 MFM-like contrast in NV-PL quenching maps
We determined for the (100)-oriented NV probe (see section 2 in SI). Figure 2(a) illustrates the relative orientation between the single NV-spin and the magnetization of the sample. We observe different contrast between the AF domains and along the FM stripes, when an off-axis external field is applied during the acquisition of NV-PL images, as can be seen in Figure 2(b) for mT. The external field is applied along , with an angle with respect to the NV-axis. The image is similar to the one obtained with MFM, but without the inherent local effect of the field from the tip; here, the external field is globally and uniformly applied to both sample and tip. High-resolution images of a FM stripe between AF domains are shown in Figure 2(c). The contrast is modified in the NV-PL quenching images when the intensity of the applied field is increased from 0 mT to 60 mT due to field compensation or competition between the external and sample stray field.
To understand the origin of the NV-PL quenching and explain the effect of the external field on the contrast of the NV-PL maps, we first performed local ODMR measurements without any external field. Figure 2(d) shows ODMR spectra measured locally when the tip is on an AF domain and on a FM stripe domain (see section 4 in SI for additional measurements). On the AF domain, we observe an ODMR spectrum with around 20 % of contrast, and the splitting is symmetric with respect to the zero-field splitting (ZFS) frequency GHz indicated by the vertical dashed line. However, in contrast, the splitting is asymmetric with respect to on the FM stripe and the contrast also decreases to less than 10% for the dip at 2.80 GHz. The features observed on the FM stripe are indications that the NV-PL quenching is a consequence of spin-mixing effects Tetienne_2012 produced by components of several mT of the sample stray magnetic field that are perpendicular to the NV-axis. At some locations along the FM stripe, the stray field can be too strong to perform any ODMR measurements (see section 4 in SI). When spin-mixing effects are present, the NV-PL intensity is reduced since there are NV-spin transitions involving states that are linear combinations of dark () and bright () states. NV-PL quenching is expected along the boundary between AF domains as these are the regions in the sample where we have FM order with a strong field.
@@
with . To simplify the description, we consider only the components of the stray field of the sample that are parallel and antiparallel to the sample surface. That is, components with and , for OOP magnetization pointing up and down, respectively. The angle between these OOP stray field components and the NV-axis are and , respectively. Due to the magnetic field compensation, different degrees of quenching are expected for regions where and are parallel or antiparallel. For example, considering the values mT for the stray field on the FM stripe (which is enough to produce spin-mixing in the case of a large-angle off-axis fieldTetienne_2012) and mT, from equation (1) and (2) we can obtain mT and mT for regions where the sample stray field components are pointing up and down, respectively. We assume that the relative degree of quenching produced by these two fields in the NV-PL signal is proportional to the difference normalized by the total field. i. e. . We interpret this as a relative contrast of between the dark and the bright areas in the NV-PL maps, leading to the MFM-like contrast in Figure 2(b). This relative contrast decreases when increasing the intensity of the external field (see section 5 in SI), being for mT. This explains the almost vanishing contrast on the FM stripe in Figure 2(c). It is noteworthy that the effect is reversible. Once the external field is switched off, a PL map similar to that for mT is recovered. For mT, the contribution is from the sample, i.e., mT, which produces the same degree of quenching. In general, is not fully 5 mT; it can be smaller or larger. However, for spin-mixing, the difference in the PL drop produced by a variation in of 1-5 mT would be around 5% Tetienne_2012, Rondin_2014. Thus, it would not be possible to differentiate between the PL quenching produced by two close values of the stray field with a significant contrast. Therefore, the contrast along the FM stripe and within the AF domains will be uniform for domains with magnetization pointing upward and downward, in the absence of an off-axis , as also observed in Figure 1(d). Additional NV-PL maps with MFM-like contrast in a different sample are shown in section 6 of the SI.

### v2

@@ Abstract
Multilayered synthetic antiferromagnets (SAFs) are artificial three-dimensional (3D) architectures engineered to create novel, complex, and stable spin textures. Non-invasive and quantitative nanoscale magnetic imaging of the two-dimensional stray field profile at the sample surface is essential for understanding the fundamental properties of the spin-structure and being able to tailor them to achieve new functionalities. However, the deterministic detection of spin textures and their quantitative characterization at the nanoscale remain challenging. Here, we use nitrogen-vacancy scanning probe microscopy (NV-SPM) under ambient conditions to perform the first quantitative vector-field magnetometry measurements in the multilayered SAF [(Co/Pt)5/Co/Ru]3/(Co/Pt)6. We investigate the static and dynamic nanoscale properties of antiferromagnetic domains with boundaries hosting "one-dimensional" ferromagnetic stripes with 100 nm of width and periodic modulation of the magnetization. By employing NV-SPM measurements in different imaging modes and involving NV-probes with various crystallographic orientations, we demonstrate distinct fingerprints emerging from GHz-range spin noise and constant stray fields on the order of several mT. This provides quantitative insights into the structure of domains and domain walls, as well as into magnetic noise associated with thermal spin-waves. Our work opens up new opportunities for quantitative vector-field magnetometry of modern magnetic materials with tailored 3D spin textures and stray field profiles, and potentially novel spin-wave dispersions—in a quantitative and non-invasive manner, with exceptional magnetic sensitivity and nanometer scale spatial resolution.
@@ 1 Introduction
Synthetic antiferromagnets (SAFs) are artificial magnetic systems that offer unique opportunities for fundamental research and applications due to their highly tunable static and dynamic properties 9, 53, 54, 56, 40. In particular, SAFs present stable magnetic textures and high-frequency spin dynamics, which are ideal properties for magnetic storage technologies and spintronic devices 30, 36, 16, 5, 33. Multilayered SAFs are systems with a more complex layered structure, designed to tune magnetic energy terms to create novel magnetic textures. 22, 21, Böhm2022, 35, 53. For instance, systems with perpendicular magnetic anisotropy (PMA) can be engineered to stabilize stripe and bubble domain states 26, to induce a domain wall (DW) transition from antiferromagnetic (AF) DWs to DWs with ferromagnetic (FM) cores20, or to stabilize FM domains separated by AF DWs when the interlayer exchange energy is reduced below the demagnetization energy 45. From a dynamic point of view, dipolar interaction can affect the spin-wave modes and induce nonreciprocity on the spin-wave dispersion 15, 24, giving rise to complex spin-wave propagation. Additionally, multilayered structures with periodic magnetic stripe domain patterns exhibit spin-wave modes and dispersion relations analogous to those of a one-dimensional magnonic crystal, which can be used to control spin-wave propagation.2, 18. Therefore, these types of multilayered architectures are a powerful and application friendly approach to develop new magnetic materials with spin textures and properties useful to improve the performance of SAFs and to design the next generation of magnetic devices 19, 60, 29, 11.
@@
The direct imaging of the two-dimensional stray magnetic field profile at the sample surface with nanoscale spatial resolution can provide insights into the properties of individual domains and DWs, which is a crucial step toward understanding and improving the fundamental properties and performance of multilayered SAFs. Magnetic force microscopy (MFM) combined with macroscopic magnetometry have been employed to investigate multilayered SAFs and to connect the nanoscale magnetism of every domain/DW with the macroscopically observed properties, such as magnetization or magnetic hysteresis 22, 20, 21, 45, Böhm2022, 28, 44, 43. However, the field of a few tens of mT of the MFM probe 1 can induce reversible changes or irreversible reorientation of the magnetic moments while scanning the sample 47, 58, 4, hindering access to the intrinsic undisturbed magnetic texture. Furthermore, macroscopic magnetometry provides only information about the average behavior of a large number of magnetic domains. Moreover, many magnetic imaging techniques probe only static properties non-quantitatively and lack the sensitivity required to investigate phenomena with weak stray magnetic field (such as AF order). Therefore, despite the enormous effort in the last few years, there is still a remaining challenge regarding the deterministic detection of magnetic textures in multilayered SAFs, along with the local, quantitative, and non-invasive characterization of both static and dynamic properties.
@@
Nitrogen-Vacancy scanning probe microscopy (NV-SPM) has recently gained immense interest as a robust quantum sensing tool. It utilizes the quantum properties of a single NV defect in diamond for quantitative and non-invasive magnetic imaging, where the magnetic vector-field sensitivity can be as high as nT with an exceptional nanoscale spatial resolution, in the range of tens of nanometers 7, 13, 42, 3, 6, 57. NV-SPM has been previously employed to investigate ultra-thin SAFs or FM thin films to determine the nature of DWs and strength of the Dzyaloshinskii-Moriya interaction (DMI) 51, 48, 17, DW motion 50, 34, spin-wave dispersion and magnetic noise 12, 14. In such systems, the magnetic thickness is typically less than 2 nm with a few mT of stray field, for which unwanted effects that often make quantitative measurements difficult, such as spin-mixing of the NV-spin states and suppression of the quantum sensor’s signal due to high off-axis stray fields, are not present. Consequently, magnetic imaging of thick multilayered SAFs with complex and novel spin textures and dynamics, producing both a large stray field of tens of mT and spin noise, remained unexplored using NV-SPM.
@@
In this work, we use NV-SPM under ambient conditions to investigate the magnetic texture of the multilayered SAF of the type [(Co/Pt)X-1/Co/Ru]N-1/(Co/Pt)X with , and . We explore the domains and DWs structure, as well as their static and dynamic properties, by performing measurements in different imaging modes and employing NV-probes with various crystallographic orientations to identify distinctive fingerprints arising from both constant stray magnetic fields and spin noise. We use optically detected magnetic resonance (ODMR) and relaxometry measurements to demonstrate the presence of both sample stray fields on the order of mT and spin noise in the GHz range. We perform qualitative magnetic imaging with commercial diamond probes containing a single NV and fabricated from (100)- and (110)-oriented diamond crystals55. By using a (111)-oriented single NV probe37, we quantitatively resolve the nanoscale DW structure and the spatial distribution of the sample stray field. Combining this with micromagnetic simulation, we confirm a three-dimensional (3D) model of the spin texture that agrees well with our quantitative measurements. Our work opens novel opportunities to understand and characterize static and dynamic properties in layered architectures with complex magnetic textures and spin-waves, quantitatively and non-invasively, with exceptionally high magnetic sensitivity and nanometer-scale spatial resolution.
@@ 2.1 NV-SPM imaging in a multilayered SAF
As illustrated in Figure 1(a), NV-SPM employs a single NV defect in a diamond scanning probe (see section 1 in the Supporting Information (SI)). The NV sensor is used to map the sample surface and spatially resolve the magnetic texture. Due to its spin-dependent photoluminescence (PL), the NV-spin state can be optically initialized with a green laser (515 nm) and read out by detecting the red NV-PL signal in the wavelength range of 650-850 nm (see methods). Qualitative magnetic imaging can be performed by mapping the intensity of the NV-PL signal while scanning the sample. For direct quantification of the sample stray field by ODMR, microwave (MW) excitation is applied to promote the NV spin transition. Thus, the stray magnetic field of the sample is determined from the Zeeman splitting produced by its component along the NV-axis, which is oriented along the crystallographic direction ⟨111⟩ in the diamond probe. The orientation of the NV-axis is defined by the angles in the laboratory frame. From our ODMR measurements under ambient conditions, we estimate a DC magnetic field sensitivity of 2 T/. The value of the detected stray field and the spatial resolution for magnetic imaging depend on the NV-sample distance, which is 36-76 nm for all our measurements (see Section 2 in SI).
@@
Figure 1(b) illustrates the structure of the multilayered SAF investigated in this work. The stacking of 6 Co layers with 5 Pt interlayers give rise to FM building blocks with PMA. i.e. out-of-plane (OOP) magnetization parallel or antiparallel to the film normal axis. The insertion of a non-magnetic material like Ru introduces an AF-coupling between adjacent FM blocks through the Ruderman–Kittel–Kasuya–Yosida exchange interaction. The strength of the AF-exchange interaction can be tuned via the Ru thickness. 9, 53, 21. In our sample, we have the AF-coupling of four FM blocks, and the full structure is described as [[Co(0.55 nm)/Pt(0.70 nm)]5/Co(0.55 nm)/Ru(0.75 nm)]3/[Co(0.55 nm)/Pt(0.70 nm)]6. The numbers in nm correspond to the thickness of every thin film along the stacking. Thus the magnetic thickness of a single FM block is nm. The hysteresis curve measured on this sample shows four discrete steps (see section 3 in SI), where every step corresponds to the magnetic reversal of one of the four FM blocks, with an average saturation magnetization kA/m for configurations with a full magnetization pointing up/down.
@@
MFM images recorded after sample saturation with an in-plane applied field reveal a remanent state constituted by one-dimensional FM stripe domains with periodic bright and dark contrast at the boundary between AF domains20, 21, as shown in Figure 1(c). The AF domain sizes are primarily in the micrometer range, with 100 nm-wide FM stripes and average periodicity of the contrast of 300 nm (see section 3 in SI) . The width of the stripes and the periodicity of the contrast can vary slightly from one stripe to another, but typically remain on the order of hundreds of nanometers. The alternating bright and dark contrast observed along the FM stripe, as well as, the slight difference in magnetic contrast between adjacent AF domains, originates from the parallel and antiparallel alignment of the sample stray field and the magnetic MFM tip. Such magnetic texture has been observed by Hellwig et al. in similar samples with various , , and layer thicknesses 20, 21. However, gaining further quantitative microscopic details on the spin texture of domains and DWs along with their dynamic properties are beyond the scope of standard MFM measurements.
@@
To this end, we employ the NV-SPM approach. Figure 1(d) shows an image acquired in the NV-PL quenching mode with a (100)-oriented NV probe. The NV-PL signal is detected pixel by pixel as the sample is scanned across the NV tip which is kept stationary at the focus of the initializing green laser beam. MW excitation is not applied during the measurements in this imaging mode. We observe large bright areas on the AF domains. Along the FM stripe domains the PL signal is quenched by about 40%. The NV-PL signal can be quenched either due to spin-mixing effects induced by strong off-axis stray fields 49 or due to a decrease of the spin-relaxation time () in case spin noise is present with frequencies resonant with the NV spin transition 12, 38. Contrary to MFM measurements, there is no periodic variation of the contrast. However, MFM-like contrast can be achieved in NV-PL maps only after applying a suitable off-axis (not aligned with the NV-axis) external magnetic field, as described in the following section.
@@ 2.2 MFM-like contrast in NV-PL quenching maps
We determined for the (100)-oriented NV probe (see section 2 in SI). Figure 2(a) illustrates the relative orientation between the single NV-spin and the magnetization of the sample. We observe a different contrast between the AF domains and along the FM stripes, when an external field off-axis is applied during the acquisition of NV-PL images, as can be seen in Figure 2(b) for mT. The external field is applied along , with an angle with respect to the NV-axis. The image is similar to the one obtained with MFM, but without the inherent local effect of the field from the tip; here, the external field is globally and uniformly applied to both the sample and the tip. High-resolution images of a FM stripe between AF domains are shown in Figure 2(c). The contrast is modified in the NV-PL quenching images when the intensity of the applied field is increased from 0 mT to 60 mT due to field compensation or competition between the external and sample stray field.
To understand the origin of the NV-PL quenching and explain the effect of the external field on the contrast of the NV-PL maps, we first performed local ODMR measurements without any external field. Figure 2(d) shows ODMR spectra measured locally when the tip is on an AF domain and on a FM stripe domain (see section 4 in SI for additional measurements). On the AF domain, we observe an ODMR spectrum with around 20 % of contrast, and the splitting is symmetric with respect to the zero-field splitting (ZFS) frequency GHz indicated by the vertical dashed line. However, in contrast, the splitting is asymmetric with respect to on the FM stripe and the contrast also decreases to less than 10% for the dip at 2.80 GHz. The features observed on the FM stripe are indications that the NV-PL quenching is a consequence of spin-mixing effects 49 produced by components of several mT of the sample stray magnetic field that are perpendicular to the NV-axis. At some locations along the FM stripe, the stray field can be too strong to perform any ODMR measurements (see section 4 in SI). When spin-mixing effects are present, the NV-PL intensity is reduced since there are NV-spin transitions involving states that are linear combinations of dark () and bright () states. NV-PL quenching is expected along the boundary between AF domains as these are the regions in the sample where we have FM order with a strong field.
@@
with . To simplify the description, we consider only the components of the stray field of the sample that are parallel and antiparallel to the sample surface. That is, components with and , for OOP magnetization pointing up and down, respectively. The angle between these OOP stray field components and the NV-axis are and , respectively. Due to the magnetic field compensation, different degrees of quenching are expected for regions where and are parallel or antiparallel. For example, considering the values mT for the stray field on the FM stripe (which is enough to produce spin-mixing in the case of a large-angle off-axis field49) and mT, from equation (1) and (2) we can obtain mT and mT for regions where the sample stray field components are pointing up and down, respectively. We assume that the relative degree of quenching produced by these two fields in the NV-PL signal is proportional to the difference normalized by the total field. i. e. . We interpret this as a relative contrast of between the dark and the bright areas in the NV-PL maps, leading to the MFM-like contrast in Figure 2(b). This relative contrast decreases when increasing the intensity of the external field (see section 5 in SI), being for mT. This explains the almost vanishing contrast on the FM stripe in Figure 2(c). It is noteworthy that the effect is reversible. Once the external field is switched off, a PL map similar to that for mT is recovered. For mT, the contribution is from the sample, i.e., mT, which produces the same degree of quenching. In general, is not fully 5 mT; it can be smaller or larger. However, for spin-mixing, the difference in the PL drop produced by a variation in of 1-5 mT would be around 5% 49, 39. Thus, it would not be possible to differentiate between the PL quenching produced by two close values of the stray field with a significant contrast. Therefore, the contrast along the FM stripe and within the AF domains will be uniform for domains with magnetization pointing upward and downward, in the absence of an off-axis , as also observed in Figure 1(d). Additional NV-PL maps with MFM-like contrast in a different sample are shown in section 6 of the SI.